OpenStudy (anonymous):

Find the solution set for the following system: 3x + 4y = 7 6x - 4y = 5

hero (hero):

Hint: Add the equations

hero (hero):

(3x + 6x) + (4y - 4y) = (7+5)

hero (hero):

9x + 0 = 12 9x = 12 x = 12/9 x = 4/3 y = 3/4 (after subbing x back in)

OpenStudy (anonymous):

then the solution set it (4/3),(3/4)??

hero (hero):

(x,y) = (4/3, 3/4)

OpenStudy (anonymous):

thank you :))